if p:q=2:3 and x:y=4:5 then value of 5px + 3qy:10px + 4qy is

Grade:

1 Answers

AskiitiansExpert Abhinav Batra
13 years ago

Dear as

p:q=2:3

let p=2k then q=3k

x:y=4:5

lt x=4u then y= 5u

5px + 3qy=5 (2k)(4u)+3(3k)(5u)=85ku

10px+4qy=10(2k)(4u)+ 4(3k)(5u)=140ku

hence 5px + 3qy:10px + 4qy = 85ku/140ku = 17/28

 

ALL the BEst

Abhinav Batra